# Policy Bifurcation in Safe Reinforcement Learning

### v2

@@ Abstract
Safe reinforcement learning (RL) offers advanced solutions to constrained optimal control problems. Existing studies in safe RL implicitly assume continuity in policy functions, where policies map states to actions in a smooth, uninterrupted manner; however, our research finds that in some scenarios, the feasible policy should be discontinuous or multi-valued, and a continuous policy can inevitably lead to constraint violations. We are the first to identify the generating mechanism of such a phenomenon, and employ topological analysis to rigorously prove the existence of policy bifurcation in safe RL, which corresponds to the contractibility of the reachable tuple. Our theorem reveals that in scenarios where the obstacle-free state space is non-simply connected, a feasible policy is required to be bifurcated, meaning its output action needs to change abruptly in response to the varying state. To train such a bifurcated policy, we propose a safe RL algorithm called multimodal policy optimization (MUPO), which utilizes a Gaussian mixture distribution as the policy output. The bifurcated behavior can be achieved by selecting the Gaussian component with the highest mixing coefficient. Experiments with vehicle control tasks show that our algorithm successfully learns the bifurcated policy and ensures satisfying safety, while a continuous policy suffers from inevitable constraint violations.
@@
Existing research in safe RL primarily focuses on integrating constrained optimization methods with RL algorithms. J. Achiam et al. [8] propose constrained policy optimization (CPO), employing trust region methods to theoretically achieve near-constraint satisfaction at each iteration. Alex Ray et al. [9] incorporate the Lagrangian multiplier method into PPO [10], resulting in PPO-Lagrangian (PPO-L), which enables constraint satisfaction and return maximization. Yu et al. [11] and Yang et al. [12] tackle the problem by exploiting feasible sets in safe RL, defining them as state space regions where policies can consistently operate without violating safety constraints. Their work enhances the understanding and application of feasible sets in safe RL.
@@
Figure 1: Limitations of continuous policies in vehicle control problems. (a), Illustration of a control problem for autonomous vehicles that encounter an obstacle and must reach a target region. (b), The trajectories of a vehicle that start from different lateral position $p_{y}$ (in meters) under the optimal policy. There is a bifurcation around $p_{y}=0$ where the vehicle steers around the obstacle in different directions. (c), Correlation between the front wheel steering angle $\delta$ (in radians) and lateral position $p_{y}$ for an autonomous vehicle at a fixed longitudinal position $p_{x}$ and speed. The dots show optimal steering responses, with jumps for varying lateral positions at $p_{y}=0$. The curve denotes a continuous policy, fails to initiate necessary avoidance maneuvers when starting from the road’s center.
@@
In terms of policy structure, existing methods apply deep neural networks to approximate the policies [8, 9, 11, 12, 13, 14]. For tasks involving continuous state and action spaces, the inherent continuity of neural networks results in policies that are also continuous, mapping system states to actions smoothly. However, existing research overlooks a serious issue: In many cases, no feasible continuous policy solution may exist for constrained OCPs. In such situations, any continuous policy learned through any safe RL methods will inevitably violate the constraints for some initial states, which is extremely detrimental in practical applications.
@@
Considering the scenario shown in Fig. 0(a), a safe RL algorithm aims to learn a policy that ensures the ego vehicle reaches the target region without colliding with the obstacle, which is in the center of the road. The ego vehicle starts initially at different positions: blue on the left, orange in the center, and green on the right side of the road. The blue vehicle should learn to detour to the left, while the green vehicle should move to the right, to avoid the obstacle. As observed, a challenge arises for the orange vehicle in the center, as shown in Fig. 0(c). If a continuous policy is adopted, the orange vehicle may execute an interpolated action that interpolates the actions taken by the blue and green vehicles, potentially leading to a direct collision with the obstacle.
@@
This issue is closely related to multimodal state-action distributions, each of which corresponds to a local optimum of the non-convex OCP. The problem of action multimodality has been noted and studied in imitation learning [15] and offline RL [16, 17, 18], where data sets are derived from different experts. Applying a continuous policy in these contexts can result in interpolation between different modes, which is often low-value and risky. This issue has received litter attention in online RL under the assumption that agents can explore and learn high-value modes while avoiding low-value interpolations. However, as we demonstrate in Fig. 0(a), a continuous policy can result in undesirable intermediate behaviors and lead to constraint violation regardless of the algorithm used. In such scenarios, feasible policies should be bifurcated, where ‘bifurcated’ signifies that the policy’s actions change abruptly in response to states. We term this phenomenon ‘policy bifurcation’. Yet, the inherent mechanisms underlying this phenomenon remain unexplored.
@@
In this work, we introduce topological concepts such as paths, loops, and homotopy, and define the critical property of contractibility. These concepts enable us to analyze constrained OCPs from a topological perspective. Our findings reveal that for constrained OCPs with non-simply connected constraints, the reachable tuple corresponding to a feasible continuous policy must be contractible. This contractibility requirement often leads to suboptimal solutions. Furthermore, when the initial state set is already noncontractible (a common occurrence), the reachable tuple cannot be contractible. This indicates that no feasible continuous policy exists, and feasible policies must be bifurcated. This finding underscores the necessity of adopting bifurcated policies in safe RL, particularly for tasks with complex constraints.
As a solution to the problem revealed by our theoretical findings, we propose the multimodal policy optimization (MUPO) algorithm. It utilizes a Gaussian mixture distribution to realize a bifurcated policy structure. Additionally, it introduces spectral normalization and forward KL divergence to facilitate the accurate learning of bifurcated policies. Our experimental results demonstrate the ability of MUPO to efficiently learn bifurcated policies that ensure safety in challenging vehicle control tasks.
@@
We focus on the constrained OCP within a continuous dynamical system, where the objective is to maximize the expected total reward while satisfying safety constraints and reaching the target set. The system is characterized by a continuous state space $\mathcal{X}$ and an action space $\mathcal{U}$, with continuous dynamics $\dot{x}(t)=f(x(t),u(t))$, where $x(t)\in\mathcal{X}$ and $u(t)\in\mathcal{U}$. The control action $u(t)$ is determined by a policy $\pi_{\theta}$ parameterized by $\theta$, such that $u(t)=\pi_{\theta}(x(t))$. Let $\mathrm{X}_{\mathrm{init}}$ be the set of initial states from which the system can start, and let $d_{\mathrm{init}}$ be the probability distribution over $\mathrm{X}_{\mathrm{init}}$. The safety constraints are denoted by the sub-zero level set of a time-invariant function $h:\mathcal{X}\rightarrow\mathbb{R}$, defining the constrained set $\mathrm{X}_{\mathrm{cstr}}=\{x\in\mathcal{X}\mid h(x)\leq 0\}$. Conversely, the region where $h(x)>0$ corresponds to the violation region $\mathrm{X}_{\mathrm{viol}}$, denoting the states that violate the safety constraints. The target state set $\mathrm{X}_{\mathrm{target}}$ denotes the desired outcomes that the system aims to achieve within the time horizon $T$. The formal statement of the constrained OCP is:
@@
Safe RL offers a promising approach to address the constrained OCP by learning a policy that improves the performance while ensuring safety. In standard safe RL, the policy $\pi_{\theta}$ is often parameterized by neural networks. Such a representation of policy usually leads to a continuous policy, which is popular in modern RL algorithms [10, 19, 20, 4]. However, the continuous policy is far from enough for the real-world applications, as it will cause unavoidable violations of safety constraints. As shown in Fig. 0(b) and Fig. 0(c), extensive scenarios require us to design a bifurcated policy to ensure safety. The mechanism of policy bifurcation for the constrained OCP remains an open problem, which is the main focus of this work.
@@ Results
Figure 2: Suboptimality of continuous policies. (a), Illustration of a constrained OCP in a 2D state space, extended with a time dimension. Due to the fact that the safety constraints are active at all times, the violation region $\mathrm{X}_{\mathrm{viol}}$ is depicted as a cylinder that extends along the time dimension. For the open-loop optimal solution, trajectories from initial states $x_{\mathrm{a}}$ and $x_{\mathrm{b}}$ to target states $x_{\mathrm{a}}^{\prime}$ and $x_{\mathrm{b}}^{\prime}$ bypass the obstacle on different sides. The loop $\ell_{aa^{\prime}b^{\prime}ba}$ cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$ without passing through $\mathrm{X}_{\mathrm{viol}}$, due to the non-simply connected property of $\mathrm{X}_{\mathrm{cstr}}$. According to our theoretical analysis, for any continuous policy acting in this manner, there must exist an initial state $x_{c}$, from which the trajectory will violate the constraints. (b), Illustration of a feasible continuous policy for the constrained OCP, showing the trajectory that avoids the obstacle from one side to ensure the feasibility of the continuous policy. Continuous policy is forced to take a significantly suboptimal trajectory compared to open-loop optimal solution, leading to a substantial loss in optimality.
@@
We introduce two key concepts in our analysis: the reachable tuple $\mathcal{R}$, and contractibility. The reachable tuple $\mathcal{R}$ denotes the set of all state-time pairs $(x,t)$ that the system can reach within a given time frame $T$, meaning the system can be in state $x$ at time $t$. Contractibility refers to the property that a set can be continuously transformed into a single point without violating the constraints of the system. It is important to note that this definition differs from the traditional notion of contractibility in topology. These concepts are crucial for evaluating the optimality and feasibility of continuous policies in constrained OCPs. Building on these concepts, we present two theorems that elucidate the limitations of continuous policies in constrained OCPs.
@@
Suboptimality of continuous policies: For a constrained OCP characterized by a Lipschitz continuous dynamic function $f$ and policy $\pi$, if the optimal solution corresponds to a reachable tuple $\mathcal{R}$ that is noncontractible, then the optimal solution cannot be achieved by continuous policy. Fig. 2 provides an illustrative example of this theorem.
@@
Infeasibility of continuous policies: If the initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible, and the target state set $\mathrm{X}_{\mathrm{target}}$ is contractible, then a constrained OCP has no feasible continuous policy. Fig. 3 provides an illustrative example of this theorem.
It is difficult to provide more specific conditions under which the optimal solution corresponds to a reachable tuple $\mathcal{R}$ that is noncontractible, and the initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible, because the desired behavior and the initial state set are often specified as part of the problem setup. However, these conditions often hold when the constrained set $\mathrm{X}_{\mathrm{cstr}}$ is non-simply connected, such as in scenarios where drones navigate around obstacles or autonomous vehicles avoid other cars.
@@
Figure 3: Infeasibility of continuous policies. This figure illustrates a constrained OCP within a 2D state space. The initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible; for instance, a closed curve $\ell_{\mathrm{init}}$ within $\mathrm{X}_{\mathrm{init}}$ encircles the violation region $\mathrm{X}_{\mathrm{viol}}$. Under a continuous policy, all points on $\ell_{\mathrm{init}}$ after the same time period form a new closed curve, such as $\ell_{\mathrm{mid}}$ in the figure. However, no continuous policy can transform $\ell_{\mathrm{init}}$ into $\ell_{\mathrm{target}}$ without intersecting $\mathrm{X}_{\mathrm{viol}}$, thus demonstrating the infeasibility of the continuous policy.
@@
Figure 4: Experimental results and visualization. (a), Bypass task illustration. (b), Encounter task illustration. (c), Open-loop control for Bypass task. (d), Open-loop control for Encounter task. (e), Training curve for the Bypass task. (f), Training curve for the Encounter task. (g), Closed-loop control effect for Bypass task. (h), Closed-loop control effect for Bypass task.
To address the limitations of continuous policies, we constructed a bifurcated policy that outputs the parameters of a Gaussian mixture distribution. By selecting the mean of the Gaussian component with the highest gate probability as the action, this approach allows for discontinuous changes in the action with respect to changes in the state. We then proposed the MUPO algorithm, developed based on maximum entropy RL [19], to train this bifurcated policy. It utilizes the forward KL divergence to enhance its ability to learn multimodal distributions, avoiding convergence to suboptimal solutions. Our experimental validation demonstrates that the bifurcated policies learned by the MUPO algorithm can effectively ensure safety and optimality in vehicle control tasks. In contrast, continuous control policies are shown to lead to constraint violations in these scenarios.
@@
We build two experimental tasks with a 3-DoF vehicle model [21]. The primary objective of these tasks is to effectively maneuver the ego vehicle along a linear trajectory while adeptly circumventing any potential obstacles. The first task, known as “Bypass,” involves a stationary obstacle on the reference path. The objective of the ego vehicle is to bypass this obstacle and then return to the reference path. The second task, known as “Encounter,” involves dealing with a vehicle approaching at a constant speed from the left side. The goal for the ego vehicle in this scenario is to maintain on the reference path while ensuring there is no collision with the approaching vehicle.
@@
The baselines for comparison are DSAC [22] and SAC [19] with continuous policies. Detailed hyperparameters can be found in the supplementary materials. We evaluate the algorithms with two metrics: maximum constraint violation and average episode return. The training curves are shown in Fig. 3(e) and Fig. 3(f). Results show that both MUPO and DSAC meet safety constraints and achieve the same high average episode returns, and outperform SAC. However, the evaluation based on random initialization cannot fully reflect the safety of the algorithms. Our next experiment will further show that only bifurcated policy can guarantee safety, while continuous policy cannot.
We plot the action variations with respect to the initial states for both the bifurcated policies of MUPO and the continuous policies of DSAC and SAC. In the Bypass task, Fig. 3(c) shows how the front-wheel steering angle $\delta$ changes with the lateral position $p_{y}$. In the Encounter task, Fig. 3(d) shows how the longitudinal acceleration $a_{x}$ changes with the longitudinal position $p_{x}$. In both tasks, the MUPO policy has learned a bifurcated policy that exhibits discontinuous transitions in actions as the state changes, which helps avoid constraint violations.
@@
Figure 5: Comparison of autonomous vehicle trajectories under continuous versus bifurcated policies in real-world execution. (a), Trajectory visualization of autonomous driving with continuous policy. (b), Trajectory visualization of autonomous driving with bifurcated policy. (c), Snapshot of autonomous vehicle executing continuous policy. (d), Snapshot of autonomous vehicle executing bifurcated policy.
@@
In our closed-loop simulation, we analyzed the behavior of the ego vehicle initialized at three specific lateral positions: +0.01 m, 0.00 m, and -0.01 m. The control curves of the front-wheel steering angle are displayed in Fig. 3(g), and the trajectories of the ego vehicle are shown in Fig. 3(h), with the obstacle represented by a black rectangle. The policy of MUPO demonstrates a bifurcation characteristic, where infinitesimal variations in the initial state induce significant shifts in behavioral modes. On the other hand, the DSAC adopts a continuous policy, adjusting behavior smoothly in response to initial state changes. However, between left and right bypass maneuvers around an obstacle, there exists no continuous transition that satisfies constraints. Consequently, employing the continuous policy inevitably leads to constraint violations. It is important to note that the initial states leading to constraint violations represent only a very small fraction of the total, which can be observed in the figures at a sub-centimeter scale. This explains why previous studies have not identified this issue: the probability of constraint violations is extremely low when initial states are randomly selected. However, the non-zero probability of such violations occurring cannot be overlooked. In scenarios where safety is critical, the adoption of bifurcated policies is essential.
@@
We explore the potential for constraint violations under continuous policies and evaluate MUPO’s capacity to ensure safety in the real-world. We utilized the Geekplus M200 mobile robot in our experiments, which was programmed to follow a linear trajectory at 0.3m/s while avoiding a centrally placed square obstacle, as illustrated in Fig. 5. We established a minimum safety distance of 0.10m. The robot’s initial positions varied, as depicted in Fig. 4(c) and Fig. 4(d). Our observations, as depicted in Fig. 4(a), confirmed that under continuous policy control, the robot could avoid the obstacle when starting from the sides. However, initiating from a central position led to a collision. In contrast, as shown in Fig. 4(b), the bifurcated policy effectively enabled the robot to avoid obstacles, regardless of the starting position. This observation is in line with our theoretical predictions. The bifurcation property of the policy is crucial in meeting constraints and approaching the global optimum.

### v3

@@ Abstract
Safe reinforcement learning (RL) offers advanced solutions to constrained optimal control problems. Existing studies in safe RL implicitly assume continuity in policy functions, where policies map states to actions in a smooth, uninterrupted manner. However, our research finds that in some scenarios, the feasible policy should be discontinuous or multi-valued, and a continuous policy can inevitably lead to constraint violations. We are the first to identify the generating mechanism of such a phenomenon, and employ topological analysis to rigorously prove the existence of policy bifurcation in safe RL, which corresponds to the contractibility of the reachable tuple. Our theorem reveals that in scenarios where the obstacle-free state space is non-simply connected, a feasible policy is required to be bifurcated, meaning its output action needs to change abruptly in response to the varying state. To train such a bifurcated policy, we propose a safe RL algorithm called multimodal policy optimization (MUPO), which utilizes a Gaussian mixture distribution as the policy output. The bifurcated behavior can be achieved by selecting the Gaussian component with the highest mixing coefficient. Experiments with vehicle control tasks show that our algorithm successfully learns the bifurcated policy and ensures satisfying safety, while a continuous policy suffers from inevitable constraint violations.
@@
Existing research in safe RL primarily focuses on integrating constrained optimization methods with RL algorithms. J. Achiam et al. [8] propose constrained policy optimization (CPO), employing trust region methods to achieve near-constraint satisfaction at each iteration theoretically. Alex Ray et al. [9] incorporate the Lagrangian multiplier method into proximal policy optimization (PPO) [10], resulting in PPO-Lagrangian (PPO-L), which enables constraint satisfaction and return maximization. Yu et al. [11] and Yang et al. [12] tackle the problem by exploiting feasible sets in safe RL, defining them as state space regions where policies can consistently operate without violating safety constraints. Their work enhances the understanding and application of feasible sets in safe RL.
@@
Figure 1: Limitations of continuous policies in vehicle control problems. (a) Illustration of a control problem for autonomous vehicles that encounter an obstacle and must reach a goal region. (b) The trajectories of a vehicle that start from different lateral positions $p_{y}$ (in meters) under the optimal policy. There is a bifurcation around $p_{y}=0$ where the vehicle steers around the obstacle in different directions. (c) Correlation between the front wheel steering angle $\delta$ (in radians) and lateral position $p_{y}$ for an autonomous vehicle starts at a fixed longitudinal position $p_{x}$ and speed. The dots show optimal steering responses, with jumps for varying lateral positions at $p_{y}=0$. The curve denotes a continuous policy that fails to initiate necessary avoidance maneuvers when starting from the road’s center.
@@
In terms of policy structure, existing methods apply deep neural networks to approximate policies [8, 9, 11, 12, 13, 14]. For tasks involving continuous state and action spaces, the inherent continuity of neural networks results in policies that are also continuous, mapping system states to actions smoothly. However, existing research overlooks a serious issue: in many cases, no feasible continuous policy solution may exist for constrained OCPs. In such situations, any continuous policy learned through any safe RL methods will inevitably violate the constraints for some initial states, which is extremely detrimental in practical applications.
@@
Considering the scenario shown in Fig. 1(a), a safe RL algorithm aims to learn a policy that ensures the ego vehicle reaches the goal without colliding with the obstacle, which is in the center of the road. The ego vehicle starts initially at different positions: blue on the left, red in the center, and green on the right side of the road. The blue vehicle should learn to detour to the left, while the green vehicle should move to the right. As observed, a challenge arises for the red vehicle in the center, as shown in Fig. 1(c). If a continuous policy is adopted, the red vehicle may perform an interpolated action based on the actions taken by the blue and green vehicles, potentially leading to a direct collision with the obstacle.
@@
This issue is closely related to multimodal state-action distributions, each of which corresponds to a local optimum of the non-convex OCP. The problem of action multimodality has been noted and studied in imitation learning [15] and offline RL [16, 17, 18], where data sets are derived from different experts. Applying a continuous policy in these contexts can result in interpolation between different modes, which is often low-value and risky. This issue has received litter attention in online RL under the assumption that agents can explore and learn high-value modes while avoiding low-value interpolations. However, as we demonstrate in Fig. 1(a), a continuous policy can result in undesirable intermediate behaviors and lead to constraint violation regardless of the algorithm used. In such scenarios, feasible policies should be bifurcated, where “bifurcated” signifies that the policy’s actions change abruptly in response to states. We term this phenomenon “policy bifurcation”. Revealing the underlying mechanisms of this phenomenon remains an unexplored area in current research.
@@
In this work, we analyze constrained OCPs from a topological perspective. We introduce topological concepts such as paths, loops, and homotopy to define the critical property of contractibility, which plays a pivotal role in our analysis. Our findings reveal that for constrained OCPs with non-simply connected constraints, the set of reachable states and times—termed the reachable tuple—corresponding to a feasible continuous policy must exhibit this contractibility. This requirement often leads to suboptimal policies. Furthermore, when the initial state set is inherently noncontractible, a common occurrence, the reachable tuple cannot be contractible, indicating the impossibility of a continuous policy to be feasible. This highlights the necessity of adopting bifurcated policies in safe RL, especially for tasks with complex safety constraints.
As a solution to the problem revealed by our theoretical findings, we propose the multimodal policy optimization (MUPO) algorithm. It utilizes a Gaussian mixture distribution to realize a bifurcated policy structure. Additionally, it incorporates spectral normalization and forward Kullback-Leibler (KL) divergence, enabling the policy to better capture multimodal action distributions, thereby facilitating the accurate learning of bifurcated policies. Our experimental results demonstrate the ability of MUPO to efficiently learn bifurcated policies that ensure safety in challenging vehicle control tasks.
@@
We focus on the constrained OCP within a continuous dynamical system, where the objective is to maximize the expected total reward while satisfying safety constraints and reaching the goal. The system is characterized by a continuous state space $\mathcal{X}$ and an action space $\mathcal{U}$, with continuous dynamics $\dot{x}(t)=f(x(t),u(t))$, where $x(t)\in\mathcal{X}$ and $u(t)\in\mathcal{U}$. The control action $u(t)$ is determined by a policy $\pi_{\theta}$ parameterized by $\theta$, such that $u(t)=\pi_{\theta}(x(t))$. Let $\mathrm{X}_{\mathrm{init}}$ be the set of initial states from which the system can start, and let $d_{\mathrm{init}}$ be the probability distribution over $\mathrm{X}_{\mathrm{init}}$. The safety constraints are denoted by the sub-zero level set of a time-invariant function $h:\mathcal{X}\rightarrow\mathbb{R}$, defining the constrained set $\mathrm{X}_{\mathrm{cstr}}=\{x\in\mathcal{X}\mid h(x)\leq 0\}$. Conversely, the region where $h(x)>0$ corresponds to the violation region $\mathrm{X}_{\mathrm{viol}}$, denoting the states that violate the safety constraints. The goal state set $\mathrm{X}_{\mathrm{goal}}$ denotes the desired outcomes that the system aims to achieve within the time horizon $T$. The formal statement of the constrained OCP is:
@@
Safe RL offers a promising approach to address the constrained OCP by learning a policy that improves performance while ensuring safety. In standard safe RL, the policy $\pi_{\theta}$ is often parameterized by neural networks. Such a representation of policy usually leads to a continuous policy, which is popular in modern RL algorithms [10, 19, 20, 4]. However, the continuous policy is far from enough for real-world applications, as it will cause unavoidable violations of safety constraints. As shown in Fig. 1(b) and Fig. 1(c), extensive scenarios require us to design a bifurcated policy to ensure safety. The mechanism of policy bifurcation for the constrained OCP remains an open problem, which is the main focus of this work.
@@ Results
Figure 2: Suboptimality of continuous policies. (a) Illustration of a constrained OCP in a 2D state space, extended with a time dimension. Due to the fact that the safety constraints are active at all times, the violation region $\mathrm{X}_{\mathrm{viol}}$ is depicted as a cylinder that extends along the time dimension. For the open-loop optimal solution, trajectories from initial states $x_{\mathrm{a}}$ and $x_{\mathrm{b}}$ to goal states $x_{\mathrm{a}}^{\prime}$ and $x_{\mathrm{b}}^{\prime}$ bypass the obstacle on different sides. The loop $\ell_{aa^{\prime}b^{\prime}ba}$ cannot be continuously contracted to a point within $\mathrm{X}_{\mathrm{cstr}}$ without passing through $\mathrm{X}_{\mathrm{viol}}$, due to the non-simply connected property of $\mathrm{X}_{\mathrm{cstr}}$. According to our theoretical analysis, for any continuous policy acting in this manner, there must exist an initial state $x_{c}$, from which the trajectory must violate the constraints. (b) Illustration of a feasible continuous policy for the constrained OCP, showing the trajectory that avoids the obstacle from one side to ensure the feasibility of the continuous policy. Continuous policy is forced to take a significantly suboptimal trajectory compared to open-loop optimal solutions, leading to a substantial loss in optimality.
@@
We introduce two key concepts in our analysis: the reachable tuple $\mathcal{R}$, and contractibility. The reachable tuple $\mathcal{R}$ denotes the set of all state-time pairs $(x,t)$ that the system can reach within a given time frame $T$, meaning the system can be in state $x$ at time $t$. Contractibility refers to the property that a set can be continuously transformed into a single point without violating the constraints of the system. It is important to note that this definition differs from the traditional notion of contractibility in topology. When the constrained set $\mathrm{X}_{\mathrm{cstr}}$ is non-simply connected, its subset could be noncontractible. This noncontractibility further leads to limitations of continuous policies in constrained OCPs, which are presented as follows:
@@
Suboptimality of continuous policies: for a constrained OCP characterized by a Lipschitz continuous dynamic function $f$, if the optimal policy corresponds to a reachable tuple $\mathcal{R}$ that is noncontractible, then the optimal policy cannot be achieved by continuous policy. Fig. 2 provides an illustrative example of this theorem.
@@
Infeasibility of continuous policies: if the initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible, and the goal state set $\mathrm{X}_{\mathrm{goal}}$ is contractible, then a constrained OCP has no feasible continuous policy. Fig. 3 provides an illustrative example of this theorem.
@@
Figure 3: Infeasibility of continuous policies. This figure illustrates a constrained OCP within a 2D state space. The initial state set $\mathrm{X}_{\mathrm{init}}$ is noncontractible; for instance, a closed curve $\ell_{\mathrm{init}}$ within $\mathrm{X}_{\mathrm{init}}$ encircles the violation region $\mathrm{X}_{\mathrm{viol}}$. Under a continuous policy, all points on $\ell_{\mathrm{init}}$ after the same time period form a new closed curve, such as $\ell_{\mathrm{mid}}$ in the figure. However, there exists no continuous deformation in topology that can transform $\ell_{\mathrm{init}}$ into $\ell_{\mathrm{goal}}$ without intersecting $\mathrm{X}_{\mathrm{viol}}$, thus demonstrating the infeasibility of the continuous policy.
@@
Figure 4: Experimental results and visualization. (a) Bypass task illustration. (b) Encounter task illustration. (c) Open-loop control for the Bypass task. (d) Open-loop control for the Encounter task. (e) Training curves for the Bypass task. (f) Training curves for the Encounter task. (g) Front-wheel steering angle $\delta$ changes over time for the Bypass task where the vehicle starts from different lateral positions $p_{y}$. (h) Trajectories for the Bypass task where the vehicle starts from different lateral positions $p_{y}$.
To address the limitations of continuous policies, we construct a bifurcated policy that outputs the parameters of a Gaussian mixture distribution. By selecting the mean of the Gaussian component with the highest gate probability as the action, this approach allows for discontinuous changes in the action with respect to changes in the state. We propose the MUPO algorithm, developed based on maximum entropy RL [19], to train this bifurcated policy. It utilizes the forward KL divergence to enhance its ability to learn multimodal distributions, avoiding convergence to suboptimal solutions. Our experimental validation demonstrates that the bifurcated policies learned by the MUPO algorithm can effectively ensure safety and optimality in vehicle control tasks. In contrast, continuous control policies are shown to lead to constraint violations in these scenarios.
@@
We build two experimental tasks with a three-degree-of-freedom vehicle model [21]. The primary objective of these tasks is to effectively maneuver the ego vehicle along a linear trajectory while adeptly circumventing any potential obstacles. The first task, referred to as “Bypass”, involves a stationary obstacle on the reference path. The objective of the ego vehicle is to bypass this obstacle and then return to the reference path. The second task, referred to as “Encounter”, involves dealing with a vehicle approaching at a constant speed from the left side. The goal for the ego vehicle in this scenario is to maintain on the reference path while ensuring there is no collision with the approaching vehicle.
@@
The baselines for comparison are distributional soft actor-critic (DSAC) [22] and soft actor-critic (SAC) [19] with continuous policies. Detailed hyperparameters can be found in the supplementary materials. First, we evaluate the algorithms with two metrics: maximum constraint violation and average episode return. The training curves are shown in Fig. 4(e) and Fig. 4(f). The results demonstrate that both MUPO and DSAC achieve comparable high average episode returns, surpassing the performance of SAC, while seemingly satisfying the safety constraints under the current evaluation setup. However, it is important to note that the evaluation based on a limited number of random initializations cannot fully reflect the safety of the learned policies. Our next experiment will further show that only bifurcated policy can guarantee safety, while continuous policy cannot.
@@
Figure 5: Comparison of autonomous vehicle trajectories under continuous versus bifurcated policies in real-world execution. (a) Trajectory visualization of autonomous driving with continuous policy. (b) Trajectory visualization of autonomous driving with bifurcated policy. (c) Snapshot of autonomous vehicle executing continuous policy. (d) Snapshot of autonomous vehicle executing bifurcated policy.
We plot the action variations against initial states for both the bifurcated policies of MUPO and the continuous policies of DSAC and SAC. In the Bypass task, Fig. 4(c) shows how the front-wheel steering angle $\delta$ changes with the initial lateral position $p_{y}$. All the policies tend to steer left (positive front-wheel angle $\delta$) when initially positioned on the left side of the road, and steer right on the right side. However, near the road center, continuous policies interpolate between left and right steering—going straight, while bifurcated policies can exhibit discontinuous transitions. In the Encounter task, Fig. 4(d) shows how the longitudinal acceleration $a_{x}$ changes with the initial longitudinal position $p_{x}$. All the policies accelerate when initially ahead, and decelerate when behind, to pass or yield, respectively. Again, only bifurcated policies exhibit discontinuous behavioral transitions. In both tasks, the MUPO learns bifurcated policies that exhibit discontinuous transitions in actions as the state changes, which helps avoid constraint violations.
@@
In our closed-loop simulation, we analyze the behavior of the ego vehicle initialized at three specific lateral positions: +0.01m, 0.00m, and -0.01m. The control curves of the front-wheel steering angle are displayed in Fig. 4(g), and the trajectories of the ego vehicle are shown in Fig. 4(h), with the obstacle represented by a black rectangle. The policy of MUPO demonstrates a bifurcation characteristic, where infinitesimal variations in the initial state induce significant shifts in behavioral modes. On the other hand, the DSAC adopts a continuous policy, adjusting behavior smoothly in response to initial state changes. However, between left and right bypass maneuvers around an obstacle, there exists no continuous transition that satisfies constraints. Consequently, employing the continuous policy inevitably leads to constraint violations. It is worth noting that the initial states leading to constraint violations constitute a minute fraction of the overall state space, as evidenced by position deviation presented at a sub-centimeter scale. This explains why previous studies have not identified this issue: the probability of constraint violations is extremely low when initial states are randomly selected. However, the non-zero probability of such violations occurring cannot be overlooked. In scenarios where safety is critical, the adoption of bifurcated policies is essential.
@@
We explore the potential for constraint violations under continuous policies and evaluate MUPO’s capacity to ensure safety in the real world. We utilize the Geekplus M200 mobile robot in our experiments, which is programmed to follow a linear trajectory at 0.3m/s while avoiding a centrally placed square obstacle, as illustrated in Fig. 5. The robot is initialized randomly from various positions, as depicted in Fig. 5(c) and Fig. 5(d). Our observations, as depicted in Fig. 5(a), confirm that the robot could avoid the obstacle when initiated from the sides under continuous policy control. However, initiating from a central position led to a collision. In contrast, as shown in Fig. 5(b), the bifurcated policy effectively facilitates obstacle avoidance for the robot, regardless of its initial position. This empirical observation aligns with our theoretical predictions, showcasing that the bifurcation property of the policy is crucial in meeting constraints and approaching the global optimum.